# Optimizing Backward Policies in GFlowNets via Trajectory Likelihood Maximization

### v1

@@ 1 Introduction
Generative Flow Networks (GFlowNets, Bengio et al., 2021) are models designed to sample compositional discrete objects, e.g. graphs, from distributions defined by unnormalized probability mass functions. They operate by constructing an object through a sequence of stochastic transitions defined by a forward policy. This policy is trained to match the marginal distribution over constructed objects with the target distribution of interest. Since this marginal distribution is generally intractable, an auxiliary backward policy is introduced, and the problem is reduced to the one of matching distributions over complete trajectories, bearing similarities with variational inference (Malkin et al., 2023).
GFlowNets have found success in various areas, such as biological sequence design (Jain et al., 2022), material discovery (Hernandez-Garcia et al., 2023), molecular optimization (Zhu et al., 2024), recommender systems (Liu et al., 2024), large language model (LLM) and diffusion model fine-tuning (Hu et al., 2023; Venkatraman et al., 2024; Uehara et al., 2024; Zhang et al., 2024), neural architecture search (Chen & Mauch, 2023), combinatorial optimization (Zhang et al., 2023), and causal discovery (Atanackovic et al., 2024).
Theoretical foundations of GFlowNets have been established in the seminal works of Bengio et al. (2021; 2023). Most of the literature has since focused on practical applications of these models, leaving their theoretical properties largely unexplored, with a few exceptions (Krichel et al., 2024; Silva et al., 2024). However, a recent line of works has brought attention to connections between GFlowNets and reinforcement learning (Tiapkin et al., 2024; Mohammadpour et al., 2024; Deleu et al., 2024; He et al., 2024a), showing that the GFlowNet learning problem is equivalent to a specific RL problem with entropy regularization (also called soft RL, Neu et al., 2017; Geist et al., 2019). This opened a new perspective for understanding GFlowNets. The importance of these findings is supported by empirical evidence, as various RL algorithms have proven useful for improving GFlowNets (Tiapkin et al., 2024; Lau et al., 2024; Morozov et al., 2024).
@@
To develop this method, we first formulate the GFlowNet training problem as a unified objective involving both forward and backward policies. We then propose an alternating minimization procedure consisting of two steps: (1) maximizing the backward policy likelihood of trajectories sampled from the forward policy and (2) optimizing the forward policy within an entropy-regularized Markov decision process that corresponds to the updated backward policy. The latter step can be achieved by any existing GFlowNet or soft RL algorithm, as it was outlined by Deleu et al. (2024). By approximating these two steps through a single stochastic gradient update, we derive an adaptive approach for combining backward policy optimization with any GFlowNet method, including soft RL methods.
@@
  • We derive the trajectory likelihood maximization (TLM) method for backward policy optimization;
@@
  • We provide extensive experimental evaluation of TLM in four tasks, confirming the findings of Mohammadpour et al. (2024), which emphasize the benefits of training the backward policy in complex environment with less structure.
@@ 2.1 GFlowNets
To formally define a generation process in GFlowNets, we introduce a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{E})$, where $\mathcal{S}$ is a state space and $\mathcal{E}\subseteq\mathcal{S}\times\mathcal{S}$ is a set of edges (or transitions). There is exactly one state, $s_{0}$, with no incoming edges, which we refer to as the initial state. All other states can be reached from $s_{0}$, and the set of terminal states with no outgoing edges coincides with the space of interest $\mathcal{X}$. Non-terminal states $s\notin\mathcal{X}$ correspond to “incomplete” objects, and edges $s\to s^{\prime}$ represent adding ”new components” to such objects, transforming $s$ into $s^{\prime}$. Let $\mathcal{T}$ denote the set of all complete trajectories $\tau=\mathopen{}\mathclose{{}\left(s_{0},s_{1},\ldots,s_{n_{\tau}}}\right)$ in the graph, where $\tau$ is a sequence of states such that $(s_{i}\to s_{i+1})\in\mathcal{E}$ and that starts at $s_{0}$ and finishes at some terminal state $s_{n_{\tau}}\in\mathcal{X}$. As a result, any complete trajectory can be viewed as a sequence of actions that constructs the object corresponding to $s_{n_{\tau}}$ starting from the ”empty object” $s_{0}$.
We say that a state $s^{\prime}$ is a child of a state $s$, if there is an edge $(s\to s^{\prime})\in\mathcal{E}$. In this case we also say that $s$ is a parent of $s^{\prime}$. Next, for any state $s$, we introduce the forward policy, denoted by $\mathcal{P}_{\mathrm{F}}(s^{\prime}|s)$ for $(s\to s^{\prime})\in\mathcal{E}$, as an arbitrary probability distribution over the set of children of the state $s$. In a similar fashion, we define the backward policy as an arbitrary probability distribution over the parents of a state $s$ and denote it as $\mathcal{P}_{\mathrm{B}}(s^{\prime}|s)$, where $(s^{\prime}\to s)\in\mathcal{E}$.
@@
where $\tau=(s_{0},s_{1},\ldots,s_{n_{\tau}})\in\mathcal{T}$. If the condition (1) is satisfied for all complete trajectories, sampling a trajectory in the forward direction using $\mathcal{P}_{\mathrm{F}}$ will result in a terminal state being sampled with probability $\mathcal{R}(x)/\mathrm{Z}$. We will call such $\mathcal{P}_{\mathrm{F}}$ a proper GFlowNet sampling policy.
@@ 2.2 GFlowNets as Soft RL
where $\lambda\geq 0$ is the regularization coefficient. Note that we use the next state instead of the more conventional action representation, as there is a one-to-one correspondence between the action taken and the resulting next state in DAG environments. Similarly, the regularized Q-values $Q^{\mathcal{P}_{\mathrm{F}}}_{\lambda}(s,s^{\prime})$ are defined as the expected discounted sum of rewards, augmented by Shannon entropy, given an initial state $s_{0}=s$ and the next state $s_{1}=s^{\prime}$. The regularized optimal policy $\mathcal{P}^{\star}_{\mathrm{F},\lambda}$ is the policy that maximizes $V^{\mathcal{P}_{\mathrm{F}}}_{\lambda}(s)$ for any state $s$.
Note. In standard RL notation, a policy is typically denoted as $\pi$. Here, we use $\mathcal{P}_{\mathrm{F}}$, consistently with GFlowNet notation, to avoid notational clutter.
@@
The main practical corollary of these results is the fact that any RL algorithm that works with entropy regularization can be utilized to train GFlowNets when $\mathcal{P}_{\mathrm{B}}$ is fixed. For example, Tiapkin et al. (2024) demonstrated the efficiency of the classical SoftDQN algorithm (Haarnoja et al., 2017) and its modified variant called MunchausenDQN (Vieillard et al., 2020). Moreover, it turns out that, under this framework, the existing GFlowNet training algorithms can be derived from existing RL algorithms. Tiapkin et al. (2024) showed that on-policy TB corresponds to policy gradient algorithms, and DB corresponds to a dueling variant of SoftDQN. At the same time, Deleu et al. (2024) showed that TB, DB and SubTB algorithms can be derived from path consistency learning (PCL, Nachum et al., 2017) under the assumption of a fixed backward policy.
@@ 2.3 Backward Policies in GFlowNets
The idea of backward policy optimization is essential to understanding the GFlowNets training procedure. In particular, the most straightforward approach used in GFlowNet literature (Malkin et al., 2022; Bengio et al., 2023) proposes to optimize the forward and backward policies directly through the same GFlowNet objective (e.g., (3)). This approach can accelerate the speed of convergence Malkin et al. (2022), at the same time potentially leading to less stable training (Zhang et al., 2022).
@@
This phenomenon motivates studying the backward policy optimization in the recent works of Mohammadpour et al. (2024) and Jang et al. (2024). Mohammadpour et al. (2024) suggested using the backward policy with maximum possible trajectory entropy, thus focusing on the exploration challenges of GFlowNets. Such policy is proven to be $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=n(s)/n(s^{\prime})$, where $n(s)$ is the number of trajectories which starts at $s_{0}$ and end at $s$. It corresponds to the uniform one if the number of paths to all parent nodes is equal. When $n(s)$ cannot be computed analytically, Mohammadpour et al. (2024) propose to learn $\log{n(s)}$, $s\in\mathcal{S}$ alongside the forward policy using its relation to the value function of the soft Bellman equation in the inverted MDP (see Definition 2 in Mohammadpour et al. (2024)).
The approach of Mohammadpour et al. (2024) showed consistently better results in less structured tasks, like QM9 molecule generation (see Section 4.3 for a detailed description). At the same time, more structured environments with a less challenging exploration counterpart show less benefits of the proposed backward training approach.
@@
In another line of work, Jang et al. (2024) claim that the existing GFlowNets training procedures tend to under-exploit the high-reward objects and propose a Pessimistic Backward Policy approach. Thus, the primary aim of Jang et al. (2024) is to focus on the exploitation of the current information about high-reward trajectories. Towards this aim, they focus on maximizing the observed backward flow $\mathsf{P}_{\mathcal{T}}^{\mathcal{P}_{\mathrm{B}}}(\tau)$ (see (2)) for trajectories leading to high-reward objects, which are stored in a replay buffer. Our approach shares similarities with this method, but differs in theoretical motivation and the choice of trajectories for $\mathcal{P}_{\mathrm{B}}$ optimization. Unfortunately, Jang et al. (2024) do not provide enough details about the choice of trajectories that are stored in their replay buffer and stochastic optimization over trajectories in the buffer. Moreover, related design choices vary between different environments, which limits reproducibility of the results.
@@
As an additional limitation of both Mohammadpour et al. (2024) and Jang et al. (2024), we mention the fact that only a single GFlowNet training objective is used in both papers (SubTB and TB respectively) to evaluate approaches for backward policy optimization, while we carry out experimental evaluation with various GFlowNet training objectives (see Section 4).
@@ 3 Trajectory Likelihood Maximization
where $\Pi_{\mathrm{F}}$ and $\Pi_{\mathrm{B}}$ represent the spaces of forward and backward policies, respectively, and $\mathsf{P}_{\mathcal{T}}^{\mathcal{P}_{\mathrm{F}}}$ and $\mathsf{P}_{\mathcal{T}}^{\mathcal{P}_{\mathrm{B}}}$ are defined in (2). It is easy to see that any solution to the problem (5) satisfies the trajectory balance constraint (1) and thus induces a proper GFlowNet sampling policy. Additionally, it is worth mentioning that for any fixed $\mathcal{P}_{\mathrm{B}}$ the problem (5) is equivalent to maximizing value
@@
During numerical experiments, we observed that enforcing stability in backward updates, particularly by using a decaying learning rate, significantly improves convergence in practice. Furthermore, as the theorem shows, this stability is essential for theoretical convergence. We discuss stability techniques we utilize alongside our algorithm in Appendix A.2.
Discussion.
We underline our approach’s similarity with a celebrated EM-algorithm (Dempster et al., 1977) and Hinton’s Wake-Sleep algorithm (Hinton et al., 1995). Both these methods also attempt to address the minimization problem using the block structure and alternating minimization approaches. Additionally, our approach can be connected to cooperative game theory. Indeed, one may interpret $\mathcal{P}_{\mathrm{F}}$ as a Forward player and $\mathcal{P}_{\mathrm{B}}$ as a Backward player, and both players attempt to minimize KL-divergence between corresponding trajectory distributions.
@@ 4 Experiments
We evaluate four GFlowNet training methods: MunchausenDQN (based on the framework of Tiapkin et al. (2024)), DB (Bengio et al., 2023), TB (Malkin et al., 2022), and SubTB (Madan et al., 2023), which we refer to as GFlowNet algorithms (denoted as $\mathcal{L}_{\mathtt{Alg}}$ in the previous section). On hypergrids, we also include results for SoftDQN (Tiapkin et al., 2024). Alongside these algorithms, we consider four strategies for learning or selecting the backward policy:
@@
  • maximum entropy backward policy (maxent(Mohammadpour et al., 2024).
@@ 4.1 Hypergrid
The environment is a $d$-dimensional hypercube with a side length equal to $H$. The state space is represented as $d$-dimensional vectors $(s_{1},\ldots,s_{d})^{\mathsf{\scriptscriptstyle T}}\in\{0,\ldots,H-1\}^{d}$ with the initial state being $(0,\ldots,0)^{\mathsf{\scriptscriptstyle T}}$. For each state $(s_{1},\ldots,s_{d-1})$, there are at most $d+1$ actions. The first action always corresponds to an exit action that transfers the state to its terminal copy, and the rest of $d$ actions correspond to incrementing one coordinate by $1$ without leaving the grid. The number of terminal states here is $|\mathcal{X}|=H^{d}$.
There are $2^{d}$ regions with high rewards near the corners of the grid, while states outside have much lower rewards. The exact expression for the rewards is given in Section A.3.2.
@@
Figure 1 presents the results. For SoftDQN, MunchausenDQN, and DB, TLM shows the fastest convergence for both “standard” and “hard” reward designs. For the SubTB algorithm, TLM shows similar performance to naive and outperforms uniform and maxent. TB is known to have difficulties in this environment (Madan et al., 2023), all approaches fail to converge under the “hard” reward design. At the same time, with the “standard” one, maxent backward shows the best convergence. An important note is that our results reproduce the findings of Tiapkin et al. (2024): for SoftDQN and MunchausenDQN training with uniform backward converges faster than with naive algorithm, while TLM shows stable improvement over uniform. The results and the ranking of algorithms are almost the same for SoftDQN and MunchausenDQN, so we leave only MunchausenDQN out of two for further experiments.

### v2

@@ 1 Introduction
Generative Flow Networks (GFlowNets, Bengio et al., 2021) are models designed to sample compositional discrete objects, e.g., graphs, from distributions defined by unnormalized probability mass functions. They operate by constructing an object through a sequence of stochastic transitions defined by a forward policy. This policy is trained to match the marginal distribution over constructed objects with the target distribution of interest. Since this marginal distribution is generally intractable, an auxiliary backward policy is introduced, and the problem is reduced to one of matching distributions over complete trajectories, bearing similarities with variational inference (Malkin et al., 2023).
GFlowNets have found success in various areas, such as biological sequence design (Jain et al., 2022), material discovery (Hernandez-Garcia et al., 2023), molecular optimization (Zhu et al., 2024), recommender systems (Liu et al., 2024), large language model (LLM) and diffusion model fine-tuning (Hu et al., 2023; Venkatraman et al., 2024; Uehara et al., 2024; Zhang et al., 2025), neural architecture search (Chen & Mauch, 2023), combinatorial optimization (Zhang et al., 2023), and causal discovery (Atanackovic et al., 2024).
Theoretical foundations of GFlowNets have been established in the seminal works of Bengio et al. (2021; 2023). Most of the literature has since focused on practical applications of these models, leaving their theoretical properties largely unexplored, with a few exceptions (Krichel et al., 2024; Silva et al., 2024). However, a recent line of works has brought attention to connections between GFlowNets and reinforcement learning (Tiapkin et al., 2024; Mohammadpour et al., 2024; Deleu et al., 2024; He et al., 2024a), showing that the GFlowNet learning problem is equivalent to a specific RL problem with entropy regularization (also called soft RL, Neu et al., 2017; Geist et al., 2019). This opened a new perspective for understanding GFlowNets. The importance of these findings is supported by empirical evidence, as various RL algorithms have proven useful for improving GFlowNets (Tiapkin et al., 2024; Morozov et al., 2024; Lau et al., 2025).
@@
We first formulate the GFlowNet training problem as a unified objective involving both forward and backward policies. We then propose an alternating minimization procedure consisting of two steps: (1) maximizing the backward policy likelihood of trajectories sampled from the forward policy and (2) optimizing the forward policy within an entropy-regularized Markov decision process that corresponds to the updated backward policy. The latter step can be achieved by any existing GFlowNet or soft RL algorithm, as it was outlined by Deleu et al. (2024). By approximating these two steps through a single stochastic gradient update, we derive an adaptive approach for combining backward policy optimization with any GFlowNet method, including soft RL methods.
@@
  • We derive the trajectory likelihood maximization (TLM) method for backward policy optimization.
@@
  • We provide an extensive experimental evaluation of TLM in four tasks, confirming the findings of Mohammadpour et al. (2024), which emphasize the benefits of training the backward policy in a complex environment with less structure.
@@ 2.1 GFlowNets
To formally define a generation process in GFlowNets, we introduce a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{E})$, where $\mathcal{S}$ is a state space and $\mathcal{E}\subseteq\mathcal{S}\times\mathcal{S}$ is a set of edges (or transitions). There is exactly one state, $s_{0}$, with no incoming edges, which we refer to as the initial state. All other states can be reached from $s_{0}$, and the set of terminal states with no outgoing edges coincides with the space of interest $\mathcal{X}$. Non-terminal states $s\notin\mathcal{X}$ correspond to “incomplete” objects and edges $s\to s^{\prime}$ represent adding ”new components” to such objects, transforming $s$ into $s^{\prime}$. Let $\mathcal{T}$ denote the set of all complete trajectories $\tau=\mathopen{}\mathclose{{}\left(s_{0},s_{1},\ldots,s_{n_{\tau}}}\right)$ in the graph, where $\tau$ is a sequence of states such that $(s_{i}\to s_{i+1})\in\mathcal{E}$ and that starts at $s_{0}$ and finishes at some terminal state $s_{n_{\tau}}\in\mathcal{X}$. As a result, any complete trajectory can be viewed as a sequence of actions that constructs the object corresponding to $s_{n_{\tau}}$ starting from the ”empty object” $s_{0}$.
We say that a state $s^{\prime}$ is a child of a state $s$ if there is an edge $(s\to s^{\prime})\in\mathcal{E}$. In this case, we also say that $s$ is a parent of $s^{\prime}$. Next, for any state $s$, we introduce the forward policy, denoted by $\mathcal{P}_{\mathrm{F}}(s^{\prime}|s)$ for $(s\to s^{\prime})\in\mathcal{E}$, as an arbitrary probability distribution over the set of children of the state $s$. In a similar fashion, we define the backward policy as an arbitrary probability distribution over the parents of a state $s$ and denote it as $\mathcal{P}_{\mathrm{B}}(s^{\prime}|s)$, where $(s^{\prime}\to s)\in\mathcal{E}$.
@@
where $\tau=(s_{0},s_{1},\ldots,s_{n_{\tau}})\in\mathcal{T}$. If the condition (1) is satisfied for all complete trajectories, sampling a trajectory in the forward direction using $\mathcal{P}_{\mathrm{F}}$ will result in a terminal state being sampled with probability $\mathcal{R}(x)/\mathrm{Z}$. We will call such $\mathcal{P}_{\mathrm{F}}$ a proper GFlowNet forward policy.
@@ 2.2 GFlowNets as Soft RL
where $\lambda\geq 0$ is the regularization coefficient. Note that we use the next state instead of the more conventional action representation, as there is a one-to-one correspondence between the action taken and the resulting next state in DAG environments. Similarly, the regularized Q-values $Q^{\mathcal{P}_{\mathrm{F}}}_{\lambda}(s,s^{\prime})$ are defined as the expected discounted sum of rewards, augmented by Shannon entropy, given an initial state $s_{0}=s$ and the next state $s_{1}=s^{\prime}$. The regularized optimal policy $\mathcal{P}^{\star}_{\mathrm{F},\lambda}$ is the policy that maximizes $V^{\mathcal{P}_{\mathrm{F}}}_{\lambda}(s)$ for any state $s$. Note: in standard RL notation, a policy is typically denoted as $\pi$. Here, we use $\mathcal{P}_{\mathrm{F}}$, consistently with GFlowNet notation, to avoid notational clutter.
@@
The main practical corollary of these results is that any RL algorithm that works with entropy regularization can be utilized to train GFlowNets when $\mathcal{P}_{\mathrm{B}}$ is fixed. For example, Tiapkin et al. (2024) demonstrated the efficiency of the classical SoftDQN algorithm (Haarnoja et al., 2017) and its modified variant called MunchausenDQN (Vieillard et al., 2020). Moreover, it turns out that, under this framework, the existing GFlowNet training algorithms can be derived from existing RL algorithms. Tiapkin et al. (2024) showed that on-policy TB corresponds to policy gradient algorithms, and DB corresponds to a dueling variant of SoftDQN. At the same time, Deleu et al. (2024) showed that TB, DB and SubTB algorithms can be derived from path consistency learning (PCL, Nachum et al., 2017) under the assumption of a fixed backward policy.
@@ 2.3 Backward Policies in GFlowNets
The idea of backward policy optimization is essential to understanding the GFlowNets training procedure. In particular, the most straightforward approach used in GFlowNet literature (Malkin et al., 2022; Bengio et al., 2023) proposes to optimize the forward and backward policies directly through the same GFlowNet objective (e.g., (3)). This approach can accelerate the speed of convergence Malkin et al. (2022), at the same time potentially leading to less stable training (Zhang et al., 2022). This phenomenon motivates studying the backward policy optimization in the recent works of Mohammadpour et al. (2024) and Jang et al. (2024).
@@
Mohammadpour et al. (2024) suggested using the backward policy with maximum possible trajectory entropy, thus focusing on the exploration challenges of GFlowNets. Such policy is proven to be $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=n(s)/n(s^{\prime})$, where $n(s)$ is the number of trajectories which starts at $s_{0}$ and end at $s$. It corresponds to the uniform one if the number of paths to all parent nodes is equal. When $n(s)$ cannot be computed analytically, Mohammadpour et al. (2024) propose to learn $\log{n(s)}$, $s\in\mathcal{S}$ alongside the forward policy using its relation to the value function of the soft Bellman equation in the inverted MDP (see Definition 2 in Mohammadpour et al. (2024)). Mohammadpour et al. (2024) utilize RL as a tool to find the maximum entropy backward policy and make the connection to RL solely for such policies. In contrast, our work theoretically considers the simultaneous optimization of the forward and the backward policy from the RL perspective, and develops an optimization algorithm grounded in it. The approach of Mohammadpour et al. (2024) showed consistently better results in less structured tasks, like QM9 molecule generation (see Section 4.3 for a detailed description). At the same time, more structured environments with a less challenging exploration counterpart show less benefits of the proposed backward training approach.
@@
In another line of work, Jang et al. (2024) claim that the existing GFlowNets training procedures tend to under-exploit the high-reward objects and propose a Pessimistic Backward Policy approach. Thus, the primary aim of Jang et al. (2024) is to focus on the exploitation of the current information about high-reward trajectories. Towards this aim, they focus on maximizing the observed backward flow $\mathsf{P}_{\mathcal{T}}^{\mathcal{P}_{\mathrm{B}}}(\tau)$ (see (2)) for trajectories leading to high-reward objects, which are stored in a replay buffer. Our approach shares similarities with this method as both compute the loss over whole trajectories but differs in theoretical motivation and the choice of trajectories for $\mathcal{P}_{\mathrm{B}}$ optimization.
@@
As a limitation of both Mohammadpour et al. (2024) and Jang et al. (2024), we mention the fact that only a single GFlowNet training objective is used in both papers (SubTB and TB respectively) to evaluate approaches for backward policy optimization, while we carry out experimental evaluation with various GFlowNet training objectives (see Section 4).
@@ 3 Trajectory Likelihood Maximization
where $\Pi_{\mathrm{F}}$ and $\Pi_{\mathrm{B}}$ represent the spaces of forward and backward policies, respectively, and $\mathsf{P}_{\mathcal{T}}^{\mathcal{P}_{\mathrm{F}}}$ and $\mathsf{P}_{\mathcal{T}}^{\mathcal{P}_{\mathrm{B}}}$ are defined in (2). It is easy to see that any solution to the problem (5) satisfies the trajectory balance constraint (1) and thus induces a proper GFlowNet forward policy. Additionally, it is worth mentioning that for any fixed $\mathcal{P}_{\mathrm{B}}$ the problem (5) is equivalent to maximizing value
@@
During numerical experiments, we observed that enforcing stability in backward updates, particularly by using a decaying learning rate and a target network, significantly improves convergence in practice. Furthermore, as the theorem shows, this stability is essential for theoretical convergence. We discuss stability techniques we utilize alongside our algorithm in Appendix A.2.
@@ 4 Experiments
We evaluate four GFlowNet training methods: MunchausenDQN (based on the framework of Tiapkin et al. (2024)), DB (Bengio et al., 2023), TB (Malkin et al., 2022), and SubTB (Madan et al., 2023), which we refer to as GFlowNet algorithms (denoted as $\mathcal{L}_{\mathtt{Alg}}$ in the previous section). On hypergrids, we also include results for SoftDQN (Tiapkin et al., 2024). Alongside these algorithms, we consider five strategies for learning or selecting the backward policy:
@@
  • maximum entropy backward policy (maxent(Mohammadpour et al., 2024);
  • pessimistic backward policy (pessimistic(Jang et al., 2024).
@@ 4.1 Hypergrid
The environment is a $d$-dimensional hypercube with a side length equal to $H$. The state space is represented as $d$-dimensional vectors $(s_{1},\ldots,s_{d})^{\mathsf{\scriptscriptstyle T}}\in\{0,\ldots,H-1\}^{d}$ with the initial state being $(0,\ldots,0)^{\mathsf{\scriptscriptstyle T}}$. For each state $(s_{1},\ldots,s_{d-1})$, there are at most $d+1$ actions. The first action always corresponds to an exit action that transfers the state to its terminal copy, and the rest of $d$ actions correspond to incrementing one coordinate by $1$ without leaving the grid. The number of terminal states here is $|\mathcal{X}|=H^{d}$. There are $2^{d}$ regions with high rewards near the corners of the grid, while states outside have much lower rewards. The exact expression for the rewards is given in Section A.3.2.
@@
Figure 1 presents the results. For SoftDQN, MunchausenDQN, and DB, TLM shows the fastest convergence for both “standard” and “hard” reward designs. For the SubTB algorithm, TLM shows similar performance to naive and outperforms uniform, maxent and pessimistic. TB is known to have difficulties in this environment (Madan et al., 2023), all approaches fail to converge under the “hard” reward design. At the same time, with the “standard” one, maxent backward shows the best convergence. An important note is that our results reproduce the findings of Tiapkin et al. (2024): for SoftDQN and MunchausenDQN training with uniform backward converges faster than with naive algorithm, while TLM shows stable improvement over uniform. The results and the ranking of algorithms are almost the same for SoftDQN and MunchausenDQN, so we leave only MunchausenDQN out of two for further experiments.
Figure 2: Top row: Bit Sequences, the number of discovered modes out of a total of 60 modes for different methods. Center row: QM9, the number of Tanimoto-separated modes with reward higher or equal to $1.125$ for different methods. Bottom row: sEH, the number of Tanimoto-separated modes with reward higher or equal to $0.875$ for different methods. Higher values indicate better performance. For each pair of a GFlowNet algorithm and a backward approach, the results are presented for the best learning rate chosen in terms of the total number of discovered modes.